The postal service of St. Vincent, an island in the West Indies, obtains a significant portion of its revenues from sales of special souvenir sheets to stamp collectors. The souvenir sheets usually contain several high-value St. Vincent stamps depicting a common theme, such as the life of Princess Diana. The souvenir sheets are designed and printed for the postal service by Imperial Printing, a stamp agency service company in the United Kingdom. The souvenir sheets cost the postal service $1.15 each. St. Vincent has been selling these souvenir sheets for $10.00 each and ordinarily sells about 61,000 units. To test the market, the postal service recently priced a new souvenir sheet at $11.00 and sales dropped to 51,000 units.

1. Calculate the contribution margin for sale price of $10.00 each or $11.00 each?

51000
$10.00 Price $11.00 Price
Unit sales 61,000 51,000
Sales $610,000 $561,000
Cost of goods sold 70,150 58,650
Contribution margin $539,850 $502,350

2. Estimate the price elasticity of demand for the souvenir sheets. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 4 decimal places.)

3. Estimate the profit-maximizing price for souvenir sheets. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Solution)

(1)

The given solution is correct. But, as contribution margin is said to be computed, it is better to write variable costs instead of the cost of goods sold.

(2)

Price Elasticity of Demand = % Change in Quantity Demanded / % Change in Price

% Change in Quantity Demanded =(51,000-61,000) x 100/ 61,000 = (16.393442623%)

% Change in Price = ($11-$10) x 100/ $10 = 10%

So, Price Elasticity of Demand = (16.393442623%)/ 10% = 1.64

(3)

The profit-maximizing price for souvenir sheets would be the selling price per souvenir sheet where the contribution margin would be maximum. Maximum contribution margin to be achieved for 51,00 units is the same as contribution achieved in the previous sales of 61,000 units.

Table showing the computation of profit-maximizing price for 51,000 souvenir sheets

Particulars Amount($)
A. Cost of goods sold (i.e Variable Cost) 58,650
B. Contribution margin 539,850
C. Sales (A+B) 598,500
D. Unit sales 51,000
E. Selling price 11.74